A spring stretches by a 100 n force. The displacement of the end of the spring increases 5.0 cm beyond its original stretches positions when the force increases 20% to 120 n. calculate the force constant of the spring and its original displacement.
1
Expert's answer
2012-11-12T11:46:35-0500
the difference in forces is proportional to difference ofdisplacement
F_2 - F_1 = k x_2 - k x_1 = k (x_2-x_1)
Hence
120-100 = k * 5cm
20 N = k* 5 cm
k = 4 N/cm